OpenStudy (anonymous):

How many solutions does the system of equations have? y=6x+2 and 3y-18x=12 One Two Infinitely many None

OpenStudy (amoodarya):

multiply first by 3 THEN ADD by the second

OpenStudy (amoodarya):

y=6x+2 y-6x=2 ---> *3

OpenStudy (anonymous):

So multiply both equations by three??

OpenStudy (amoodarya):

no first one

OpenStudy (amoodarya):

y-6x=2 ---> *3 goes to 3y-18x=6 now compare with other one 3y-18x=12 that means 12=6 now wahts the answer ?

OpenStudy (anonymous):

The lines are parallel, no solution.

OpenStudy (anonymous):

Ohh.. No solution. Thank you!

OpenStudy (anonymous):

What about x-2y=6 and 3x-6y=18 ???

OpenStudy (superdavesuper):

multiply 1st eqn by 3 n u got the same eqn as the 2nd one. so theres infinite no. of solution as well :)

OpenStudy (anonymous):

wouldn't there be one?

OpenStudy (superdavesuper):

Nope u have 2 unknown, x and y, but only one eqn cuz' the 1st n 2nd eqn are the same.

OpenStudy (anonymous):

so, Infinitely many?

OpenStudy (superdavesuper):

Correct :)

OpenStudy (anonymous):

Thanks. :D